40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 20 01:40:37 2013, Manasa, Update, Alignment, Final IFO alignment- in progress 
    Reply  Wed Feb 20 03:13:10 2013, yuta, Update, Alignment, clipping centering checklist ClippingCenteringChecklist.pdf
       Reply  Wed Feb 20 20:58:37 2013, yuta, Update, Alignment, clipping centering checklist ClippingCenteringChecklist.pdf
          Reply  Wed Feb 20 21:12:37 2013, rana, Update, Alignment, clipping centering checklist 
             Reply  Wed Feb 20 21:56:08 2013, yuta, Update, Alignment, clipping centering checklist 
    Reply  Wed Feb 20 10:13:41 2013, yuta, Update, Alignment, PRC flashing brighter than last week 
Message ID: 8114     Entry time: Wed Feb 20 03:13:10 2013     In reply to: 8113     Reply to this: 8122
Author: yuta 
Type: Update 
Category: Alignment 
Subject: clipping centering checklist 

I attached clipping/centering checklist for the alignment.
Blue ones are the ones we checked today. Red ones should be checked tomorrow. Circles indicate centering on the optics, rectangles indicate clipping check, and arrows indicate retro-reflecting or bounces.
We found mis-centering on MMT1, PR2 and SR3 tonight (by ~0.5 beam diameter). They are also indicated.

I think we don't want to touch MMT1 and PR2 anymore, because they change input beam pointing.
I'm a little bit concerned about high beam on SR3, because we had PRC flashing in vertical higher order modes. We also see ETMX slider values high in pitch (~ 5.4).
Also, the diameter of ETMX reflected beam on ITMX looked larger and dimmer than ITMX transmitted beam, which doesn't seem reasonable.


Wednesday, Feb 20:
 - tweak TT1/TT2 and PRM so PRC flashes
 - re-check Yarm/Xarm bounces
 - center beam on all AS optics, starting from SR2
 - make sure REFL and AS is clear
 - check if TRY/TRX are coming out from the ends
 - check beam centering on mirrors in IMC/OMC chamber as far as you can reach
 - inject green from both ends
 - make sure green beams are not clipped by mirrors on BS chamber, IMC/OMC chamber
 - re-center all oplevs, with no clipping
 - check all OSEM values
 - take pictures of flipped PR2 and input TTs (and everything)
 - close all heavy doors and put the access connector back

Thursday, Feb 21:
 - make sure we can lock PRMI
 - start pumping down when Steve arrives

Attachment 1: ClippingCenteringChecklist.pdf  1.059 MB  | Hide | Hide all
ClippingCenteringChecklist.pdf ClippingCenteringChecklist.pdf ClippingCenteringChecklist.pdf ClippingCenteringChecklist.pdf ClippingCenteringChecklist.pdf ClippingCenteringChecklist.pdf ClippingCenteringChecklist.pdf
ELOG V3.1.3-